Part A: A cylinder, with a piston pressing down with a constant pressure, is filled with...

Part A:

A cylinder, with a piston pressing down with a constant pressure, is filled with 2.00 moles of a gas (n1), and its volume is 49.0 L (V1). If 0.600 mole of gas leak out, and the pressure and temperature remain the same, what is the final volume of the gas inside the cylinder? Express your answer with the appropriate units.

Part B:

A sample of gas in a cylinder as in the example in Part A has an initial volume of 54.0 L , and you have determined that it contains 1.10 moles of gas. The next day you notice that some of the gas has leaked out. The pressure and temperature remain the same, but the volume has changed to 13.5 L . How many moles of gas (n2) remain in the cylinder?

Express your answer with the appropriate units.
Answer #1

Part A)

Given,

Number of moles of gas (n1) = 2.00 mol of gas

Volume of the gas(V1) = 49.0 L

If 0.600 mol of gas leaks out,

The number of moles of gas remains(n2) = 2.00 mol - 0.600 mol = 1.40 mol

Now, We know, the Avogadro's law at two different sets of conditions,

V1 / n1 = V2 / n2

Rearranging the formula,

V2 = n2 x V1 / n1

Substituting the known values,

V2 = (1.40 mol x 49.0 L) / 2.00 mol

V2 = 34.3 L

Thus, the final volume of the gas inside the cylinder is 34.3 L

Part B)

Given,

Number of moles of gas (n1) = 1.10 mol of gas

Volume of the gas(V1) = 54.0 L

Volume of the gas(V2) = 13.5 L

Now, We know, the Avogadro's law at two different sets of conditions,

V1 / n1 = V2 / n2

Rearranging the formula,

n2 = V2 x n1 / V1

Substituting the known values,

n2 = 13.5 L x 1.10 mol / 54.0 L

n2 = 0.275 mol

Thus, the number of moles of gas remains in the cylinder is 0.275 mol